OpenStudy (anonymous):

Find the partial derivative of f(x,y)=(3x+3y)e^y

OpenStudy (anonymous):

y'[x] == (-3 E^y + Derivative[1, 0][f][x, y])/(3 E^y + 3 E^y x + 3 E^y y - Derivative[0, 1][f][x, y])

OpenStudy (zarkon):

\[f_x=3e^y\] \[f_y=(3x+3y)e^y+3e^y\]

OpenStudy (anonymous):

f'x = 3e^y(3x+3y) f'y = 3e^y(3x+3y)

OpenStudy (anonymous):

why dn't you just multpply them out... i beleive that zarkon is correct

OpenStudy (zarkon):

i am ;)

OpenStudy (anonymous):

outside inside rule e^y will stay e^y always, leave inside then multiply inside respect to x and y.

OpenStudy (anonymous):

zarkon is probably right though go with zarkon.

OpenStudy (anonymous):

if you multiply it out Goat you get 3xe^y+3ye^y fx =3e^y + 0 fy=3xe^y+3e^y=3e^y(x+1)

OpenStudy (anonymous):

could be errors but i just did that without paper

OpenStudy (zarkon):

use the product rule for fy

OpenStudy (anonymous):

ahh yes both y's so yeah

OpenStudy (anonymous):

but the first is correct

OpenStudy (anonymous):

yep i see thanks i am working on my understanding of this as well.

OpenStudy (anonymous):

so how would I find fxx and fyy

OpenStudy (zarkon):

just take the derivative again

OpenStudy (zarkon):

\[f_{xx}=0\] \[f_{yy}=(3y+3x+6)e^y\]

OpenStudy (anonymous):

and fxy = 3e^y?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thank you!